### v2

@@ 2 $s_{2}$ $o_{2}$ $n_{2}$ $o_{2}/n_{2}$
The difference between the two schedules lies in the cross terms: $o_{1}n_{2}$ versus $o_{2}n_{1}$. Therefore, if $o_{1}/n_{1}<o_{2}/n_{2}$, then $o_{1}n_{2}<o_{2}n_{1}$, so prioritizing batch 1 minimizes TEL. Otherwise, prioritizing batch 2 minimizes TEL. Crucially, this optimal decision rule precisely aligns with selecting the batch with smaller $F(\mathcal{X})$, since $F(\mathcal{X})=o_{i}/n_{i}$ for each batch $i$. This demonstrates that the F-metric naturally guides the scheduler to the optimal choice.
Examples 3.1 and 3.2 demonstrate that the F-metric effectively guides scheduling decisions under different workloads. Beyond these illustrative cases, the metric exhibits valuable structural properties that contribute to robust performance in complex scenarios with varying input sizes ($s_{i}$) and output lengths ($o_{i}$).
A key challenge arises from this variability in output sizes: when requests are batched together, they do not necessarily finish processing at the same time. Under our model, once a request completes, its memory resources are freed, allowing new requests to be dynamically added to the batch. However, if output sizes in a batch vary too much, requests finish asynchronously, forcing the scheduler to switch from batch selection to incremental request insertion.
To address this, one such property—formalized in Lemma 3.3—ensures that within any selected batch, no single request’s output length dominates the average. This balancing characteristic prevents straggler effects and helps maintain predictable processing times, while also providing the analytical leverage needed for our competitive ratio guarantee in Section 4.
@@
For any batch $\mathcal{X}$ selected by the algorithm, let $o_{\max}=\max_{r_{i}\in\mathcal{X}}o_{i}$ and $\bar{o}=\frac{\sum_{r_{i}\in\mathcal{X}}o_{i}}{|\mathcal{X}|}$. Then,
@@
Proof of Lemma 3.3: From the algorithm’s selection criterion, any batch $\mathcal{X}$ must satisfy
@@
Lemma 3.3 ensures that no single request’s output length dominates the batch average, making processing times predictable even under significant workload heterogeneity.
After having introduced the intuition and property behind the F-metric, we start to describe our scheduling algorithm, Sorted-F. The algorithm itself operates in two sequential phases, executed over discrete timesteps $t$. It tracks three request sets: pending requests $\mathcal{R}^{(t)}$, active requests $\mathcal{V}^{(t)}$, and newly initiated requests $\mathcal{U}^{(t)}$. The memory consumption at any future timestep $t^{*}$ is calculated using
$M(\mathcal{X},t^{*})=\sum_{r_{i}\in\mathcal{X}}\left(s_{i}+t^{*}-p_{i}\right)\cdot\mathds{1}_{\{o_{i}\geq t^{*}-p_{i}\}},$
which accounts for the cumulative growth of KV cache entries as tokens are generated. The complete procedure is formalized in Algorithm 1.
@@
3:while $\mathcal{I}$ do
4:  $\mathcal{X}^{*}\leftarrow\arg\min_{\mathcal{X}\subseteq\mathcal{I}}(F(\mathcal{X}),-|\mathcal{X}|)$ subject to $M(\mathcal{X},p_{i}+o_{i})\leq M,\forall r_{i}\in\mathcal{X}$
5:  Sort $r_{i}\in\mathcal{X}^{*}$ in ascending order of $o_{i}$
6:  $\mathcal{I}^{\prime}\leftarrow\mathcal{I}^{\prime}+\mathcal{X}$, $\mathcal{I}\leftarrow\mathcal{I}-\mathcal{X}$
7:end while
8:Phase 2: Scheduling Execution
@@
10:while $\mathcal{R}^{(t)}+\mathcal{V}^{(t)}$ do
11:  $t\leftarrow t+1$, $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t-1)}-\mathcal{U}^{(t-1)}$, $\mathcal{V}^{(t)}\leftarrow\mathcal{V}^{(t-1)}+\mathcal{U}^{(t-1)}$, $\mathcal{U}^{(t)}\leftarrow[]$
12:  for $r_{i}\in\mathcal{R}^{(t)}$ in sorted order do
13:   if $M(\mathcal{V}^{(t)}+\mathcal{U}^{(t)},p_{j}+o_{j})\leq M,\forall r_{j}\in\mathcal{V}^{(t)}+\mathcal{U}^{(t)}+[r_{i}]$ then
14:     $\mathcal{U}^{(t)}\leftarrow\mathcal{U}^{(t)}+[i]$
15:   else
16:     Break
17:   end if
18:  end for
@@
20:end while
In Phase 1 (line 1-7), Sorted-F iteratively constructs request batches $\mathcal{X}^{*}$ that minimize $F(\mathcal{X})$ while respecting memory constraints $M(\mathcal{X},p_{i}+o_{i})\leq M$. Selected requests in a same batch are sorted by ascending $o_{i}$. The sorting criterion ensures strict ordering between batches: the request with maximal output length $o_{i}^{\max}$ in batch $B_{k}$ immediately precedes the request with minimal output length $o_{i}^{\min}$ in batch $B_{k+1}$. This generates a sorted request set $\mathcal{I}^{\prime}$.
Phase 2 (line 8-20) handles real-time token generation by: dynamically updating request sets at each timestep; admitting new requests in $o_{i}$-sorted order only if they preserve memory constraints at all future timesteps; and processing one token per active request concurrently.
The algorithm’s design provides significant advantages. First, the output length bound $o_{i}\leq 2\cdot\bar{o}$ ensures predictable batch processing times. Second, the F-metric prevents both small-$s_{i}$ large-$o_{i}$ requests from monopolizing early resources and large-$s_{i}$ small-$o_{i}$ requests from creating fragmentation. Third, the two-phase structure enables efficient hardware implementation by decoupling optimization from execution. Through these mechanisms, our approach effectively addresses the scheduling pathology from first-come-first-serve or shortest-first demonstrated in Theorem 2.1.
@@ 4 Proof of Constant Competitive Ratio
To prove Theorem 4.1, we transform Sorted-F sequentially: $\textup{{Sorted-F}}\to\textup{{Sorted-F}}_{\mathrm{separate}}$ (Theorem 4.2) $\to\textup{{Sorted-F}}_{\mathrm{group}}$ (Theorem 4.5) $\to\textup{{Sorted-F}}_{\mathrm{align}}$ (Theorem 4.8). At each step, we deliberately modify the schedule’s structure, but crucially, we prove that the cumulative latency increase remains within constant factors. In parallel, we restructure Optimal into $\textup{{Optimal}}_{\mathrm{align}}$ (Theorem 4.10). The proof culminates in a direct comparison between $\textup{{Sorted-F}}_{\mathrm{align}}$ and $\textup{{Optimal}}_{\mathrm{align}}$ (Theorem 4.12).
Before delving into the technical steps, we clarify the asymptotic regime in which the analysis is carried out. Throughout the proof, we consider a large-instance setting where the request set $\mathcal{I}$ is sufficiently large relative to the memory budget $M$. Recall from Section 2 that each request has size $s_{i},o_{i}=\epsilon(M)$, meaning that individual requests consume a vanishing fraction of memory. W.L.O.G, we further assume that the total number of requests $n=n(M)$ satisfies $\frac{n(M)\,\epsilon(M)}{M}\;\longrightarrow\;\infty$ as $M\to\infty$. Otherwise, by processing the finite amount of requests, any scheduling algorithm can have a constant competitive ratio. In words, the request set is effectively infinite, and its cardinality grows much faster than the maximum number of requests that can fit into a single batch of capacity $M$. This ensures that the contribution of any fixed group becomes negligible compared with the “tail” of remaining requests, thereby justifying the asymptotic bounds used in later arguments.
With this standing assumption in place, we now begin the first step of the proof. Let $\{\mathcal{X}_{k}\}$ denote the sequence of batches iteratively generated in Phase 1 of Sorted-F. To construct $\textup{{Sorted-F}}_{\mathrm{separate}}$, we process $\{\mathcal{X}_{k}\}$ one at a time. Specifically, for any batch $\mathcal{X}_{k}$, all requests in it start simultaneously, and the subsequent batch $\mathcal{X}_{k+1}$ does not begin until all requests in $\mathcal{X}_{k}$ completed. This sequential batch processing simulates a worst-case scenario by ensuring that no request from subsequent batches is initiated until the preceding batch has fully finished.
Let the requests in $\mathcal{X}_{k}$, ordered by increasing response length, be denoted as $r_{k,1},r_{k,2},\ldots,r_{k,n_{k}}$, where $n_{k}=|\mathcal{X}_{k}|$. Then the process of transforming Sorted-F into $\textup{{Sorted-F}}_{\mathrm{separate}}$ is depicted in Figure 1.
Figure 1: From Sorted-F to $\textup{{Sorted-F}}_{\mathrm{separate}}$
@@
For any $\mathcal{I}$, the total end-to-end latency of Sorted-F is bounded above by that of $\text{{Sorted-F}}_{\text{separate}}$, i.e.,
@@
Proof 4.3
Proof of Theorem 4.2: Let $p_{k,j}$ and $p^{\prime}_{k,j}$ denote the initiation times of request $r_{k,j}$ under Sorted-F and $\textup{{Sorted-F}}_{\mathrm{separate}}$, respectively. To prove the theorem, we show that for all batches $\mathcal{X}_{k}$ and all $j\in\{1,\dots,n_{k}\}$,
$p_{k,j}\leq p^{\prime}_{k,j}.$
This inequality holds because $\textup{{Sorted-F}}_{\mathrm{separate}}$ enforces sequential batch processing: no request in $\mathcal{X}_{k+1}$ starts until all requests in $\mathcal{X}_{k}$ complete. Thus, $\textup{{Sorted-F}}_{\mathrm{separate}}$ delays requests relative to Sorted-F, which allows overlapping batch execution.
To formalize this, we analyze the memory usage constraints. Consider modifying $\textup{{Sorted-F}}_{\mathrm{separate}}$ by iteratively advancing the initiation times $p^{\prime}_{k,j}$ to their earliest feasible points, while respecting memory limits. For each $\mathcal{X}_{k}$, process requests in order $j=1,2,\cdots,n_{k}$. After advancing $r_{k,j-1}$ ($j\geq 2$), the maximum memory usage in the interval $\max\{p^{\prime}_{k,1},p_{k,j-1}+o_{k,j-1}\},p^{\prime}_{k,j}+o_{k,j}]$ is
$\sum_{i=j}^{n_{k}}s_{i}+o_{k,n_{k}}\cdot(n_{k}-j+1).$
Advancing $r_{k,j}$ reduces this usage to
$\sum_{i=j}^{n_{k}}s_{i}+(o_{k,n_{k}}-(p^{\prime}_{k,j}-p_{k,j}))^{+}\cdot(n_{k}-j+1)\leq\sum_{i=j}^{n_{k}}s_{i}+o_{k,n_{k}}\cdot(n_{k}-j+1).$
@@
$\sum_{r_{i}\in\mathcal{X}_{k}}(s_{i}+o_{i})>M-\epsilon(M),$
where $\epsilon(M)$ is a negligible term asymptotically smaller than $M$.
Given Definition 4.4, we present Theorem 4.5 to establish an upper bound of the total end-to-end latency of $\textup{{Sorted-F}}_{\mathrm{separate}}$.
@@
Proof 4.6
Proof of Theorem 4.5: We construct aggregated groups $\{\mathcal{Y}_{m}\}$ via
@@
where $b_{m}=\sum_{j=1}^{m}a_{j}$ tracks cumulative batches, $a_{j}$ counts batches in $\mathcal{Y}_{j}$, and the addition represents concurrent initiation of all requests in constituent batches. Figure 2 illustrates this transformation from $\textup{{Sorted-F}}_{\mathrm{separate}}$ to $\textup{{Sorted-F}}_{\mathrm{group}}$.
Figure 2: From $\textup{{Sorted-F}}_{\mathrm{separate}}$ to $\textup{{Sorted-F}}_{\mathrm{group}}$
Each group $\mathcal{Y}_{m}$ is constructed such that $b_{m}$ is the smallest integer $k>b_{m-1}$ satisfying the inequality
@@
Next, we present Lemma 4.7 to characterize the geometric relationship between processing times of batches within each group, where the processing time of a batch is defined as the longest response length among all its requests.
@@
The proof of Lemma 4.7 appears in Appendix 9. By Lemma 4.7, we can derive
$\sum_{i=1}^{a_{m}}o_{b_{m-1}+i,n_{b_{m-1}+i}}$ $\leq\sum_{i=1}^{a_{m}}\left(\frac{1}{2}\right)^{\lfloor\frac{a_{m}-i}{2}\rfloor}\cdot o_{b_{m},n_{b_{m}}}$
$\leq 2\cdot\frac{1-\left(\frac{1}{2}\right)^{\lfloor\frac{a_{m}-1}{2}\rfloor}}{1-\frac{1}{2}}\cdot o_{b_{m},n_{b_{m}}}$
$\leq 4\cdot o_{b_{m},n_{b_{m}}}.$
Therefore, given any $\mathcal{I}$,
$=\frac{1}{4}\cdot\sum_{m=1}^{\infty}\left(\sum_{l=b_{m-1}+1}^{b_{m}}o_{l,n_{l}}\cdot\left(\sum_{k=b_{m}+1}^{\infty}n_{k}+\epsilon\left(\sum_{k=b_{m}+1}^{\infty}n_{k}\right)\right)\right)+\sum_{k=1}^{\infty}\sum_{i=1}^{n_{k}}o_{k,i}$
$=\frac{1}{4}\cdot\sum_{m=1}^{\infty}\left(\sum_{l=b_{m-1}+1}^{b_{m}}o_{l,n_{l}}\cdot\left(\sum_{k=b_{m}+1}^{\infty}n_{k}+\sum_{k=b_{m-1}+1}^{m}n_{k}\right)\right)+\sum_{k=1}^{\infty}\sum_{i=1}^{n_{k}}o_{k,i}$
@@
Next, we focus on the terminal batches $\{\mathcal{X}_{b_{m}}\}$. To construct $\text{{Sorted-F}}_{\text{align}}$, we replace each request’s original response length in $\mathcal{X}_{b_{m}}$ with the batch’s average response length:
$\bar{o}_{b_{m}}=\frac{\sum_{i=1}^{n_{b_{m}}}o_{b_{m},i}}{n_{b_{m}}}.$
This conversion from $\textup{{Sorted-F}}_{\mathrm{group}}$ to $\textup{{Sorted-F}}_{\mathrm{align}}$ is shown in Figure 3.
Figure 3: From $\textup{{Sorted-F}}_{\mathrm{group}}$ to $\textup{{Sorted-F}}_{\mathrm{align}}$
@@
$\eta_{m}$ $\geq\frac{\sum_{r_{i}\in\mathcal{Y}_{m}}(2s_{i}+o_{i})}{2\cdot M}>\frac{M-\epsilon(M)}{2\cdot M}=\frac{1}{2}.$ (2)

### v3

@@ 2 $s_{2}$ $o_{2}$ $n_{2}$ $o_{2}/n_{2}$
The difference between the two schedules lies in the cross terms: $o_{1}n_{2}$ versus $o_{2}n_{1}$. Therefore, if $o_{1}/n_{1}<o_{2}/n_{2}$, then $o_{1}n_{2}<o_{2}n_{1}$, so prioritizing batch 1 minimizes TEL. Otherwise, prioritizing batch 2 minimizes TEL. This decision rule aligns with selecting the batch with smaller $F(\mathcal{X})$, since $F(\mathcal{X})=o_{i}/n_{i}$ for each batch $i$.
Examples 3.1 and 3.2 illustrate how the F-metric balances batch size and response lengths in scheduling decisions.
A key challenge arises from variability in output sizes: when requests are batched together, they do not necessarily finish processing at the same time. Under our model, once a request completes, its memory resources are freed, allowing new requests to be dynamically added. If output sizes in a batch vary too much, requests finish asynchronously, forcing the scheduler to switch from batch selection to incremental request insertion.
To address this, one useful property—formalized in Lemma 3.3—ensures that within any selected batch, no single request’s output length dominates the average. This balancing characteristic helps avoid stragglers and also provides analytical leverage in Section 4.
@@
For any batch $\mathcal{X}$ selected in Phase 1, let $o_{\max}=\max_{r_{i}\in\mathcal{X}}o_{i}$ and $\bar{o}=\frac{\sum_{r_{i}\in\mathcal{X}}o_{i}}{|\mathcal{X}|}$. Then,
@@
Proof of Lemma 3.3: Let $r_{i_{\max}}\in\arg\max_{r_{i}\in\mathcal{X}}o_{i}$ and define $\mathcal{X}^{-}:=\mathcal{X}\setminus\{r_{i_{\max}}\}$. Since $\mathcal{X}^{-}$ is a feasible subset whenever $\mathcal{X}$ is feasible, the Phase 1 selection rule, minimizing $F(\cdot)$ over feasible subsets, implies
$F(\mathcal{X}^{-})\geq F(\mathcal{X}),$
i.e.,
@@
After introducing the intuition and a useful property behind the F-metric, we now describe our scheduling algorithm, Sorted-F. Let $\mathcal{I}$ denote the set of all requests in the instance. The algorithm has two phases. Phase 1 constructs an ordered list $\mathcal{I}^{\prime}$ by repeatedly selecting a feasible batch with the smallest $F(\cdot)$, with ties broken by preferring a larger batch. Phase 2 then executes the schedule over discrete timesteps $t=1,2,\ldots$ by maintaining: pending requests $\mathcal{R}^{(t)}$ (not yet started), active requests $\mathcal{V}^{(t)}$ (started but not yet completed), and newly admitted requests $\mathcal{U}^{(t)}$ (admitted at time $t$ and starting at time $t$). Each request $r_{i}$ has a unique start time $p_{i}$. The memory consumption at a future timestep $t^{*}$ is
$M(\mathcal{X},t^{*})=\sum_{r_{i}\in\mathcal{X}}\left(s_{i}+t^{*}-p_{i}\right)\cdot\mathds{1}_{\{o_{i}\geq t^{*}-p_{i}\}}.$
The complete procedure is formalized in Algorithm 1.
@@
3:while $\mathcal{I}\neq\emptyset$ do
4:  $\mathcal{X}^{*}\leftarrow\arg\min_{\mathcal{X}\subseteq\mathcal{I}}(F(\mathcal{X}),-|\mathcal{X}|)$
5:    subject to $M(\mathcal{X},o_{j})\leq M,\ \forall r_{j}\in\mathcal{X}$
6:  Sort requests in $\mathcal{X}^{*}$ in nondecreasing order of $o_{i}$
7:  Append the ordered list $\mathcal{X}^{*}$ to the end of $\mathcal{I}^{\prime}$
8:  $\mathcal{I}\leftarrow\mathcal{I}\setminus\mathcal{X}^{*}$
9:end while
10:Phase 2: Scheduling Execution
@@
12:while $\mathcal{R}^{(t)}\neq\emptyset\ \textbf{or}\ \mathcal{V}^{(t)}\neq\emptyset$ do
13:  $t\leftarrow t+1$
14:  $\mathcal{U}^{(t)}\leftarrow\emptyset$
15:  $\mathcal{V}^{(t)}\leftarrow\{r_{j}\in\mathcal{V}^{(t-1)}:p_{j}+o_{j}\geq t\}$
16:  for $r_{i}\in\mathcal{R}^{(t-1)}$ in order do
17:   tentatively set $p_{i}\leftarrow t$
18:   $\widehat{\mathcal{U}}\leftarrow\mathcal{U}^{(t)}\cup\{r_{i}\}$
19:   if $M(\mathcal{V}^{(t)}\cup\widehat{\mathcal{U}},\,p_{j}+o_{j})\leq M,\ \forall r_{j}\in\mathcal{V}^{(t)}\cup\widehat{\mathcal{U}}$ then
20:     $\mathcal{U}^{(t)}\leftarrow\widehat{\mathcal{U}}$
21:   else
22:     break
23:   end if
24:  end for
25:  $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t-1)}\setminus\mathcal{U}^{(t)}$
26:  $\mathcal{V}^{(t)}\leftarrow\mathcal{V}^{(t)}\cup\mathcal{U}^{(t)}$
@@
28:end while
In Phase 1, Sorted-F iteratively selects batches $\mathcal{X}^{*}$ that minimize $(F(\mathcal{X}),-|\mathcal{X}|)$ over feasible subsets, i.e., it first minimizes $F(\mathcal{X})$ and breaks ties by preferring a larger batch. The selected requests within each batch are then sorted by nondecreasing $o_{i}$, and the resulting ordered batches are concatenated into the ordered list $\mathcal{I}^{\prime}$.
In Phase 2, the algorithm executes over timesteps by (i) removing completed requests, (ii) admitting new requests from $\mathcal{R}^{(t-1)}$ in the order induced by $\mathcal{I}^{\prime}$ as long as the memory feasibility checks remain satisfied, and (iii) generating one token per active request concurrently.
@@ 4 Proof of Constant Competitive Ratio
To prove Theorem 4.1, we transform Sorted-F sequentially: $\textup{{Sorted-F}}\to\textup{{Sorted-F}}_{\mathrm{separate}}$ (Theorem 4.4) $\to\textup{{Sorted-F}}_{\mathrm{group}}$ (Theorem 4.7) $\to\textup{{Sorted-F}}_{\mathrm{align}}$ (Theorem 4.13). At each step, we deliberately modify the schedule’s structure, but crucially, we prove that the cumulative latency increase remains within constant factors. In parallel, we restructure Optimal into $\textup{{Optimal}}_{\mathrm{align}}$ (Theorem 4.15). The proof culminates in a direct comparison between $\textup{{Sorted-F}}_{\mathrm{align}}$ and $\textup{{Optimal}}_{\mathrm{align}}$ (Theorem 4.19). For clarity, we state each transformation as a self-contained mathematical definition (Definitions 4.34.16). Figures 25 are schematic illustrations of these definitions and are not used as substitutes for the formal constructions.
Remark 4.2 (Intermediate schedules and feasibility)
All schedules introduced in the proof as transformations of Sorted-F or Optimal (e.g., $\textup{{Sorted-F}}_{\mathrm{separate}}$, $\textup{{Sorted-F}}_{\mathrm{group}}$, $\textup{{Sorted-F}}_{\mathrm{align}}$, and $\textup{{Optimal}}_{\mathrm{align}}$) are analytical intermediates used only to compare TEL values. They are not meant to represent executable schedules and may violate the memory constraint. Whenever the budget $M$ appears, it is invoked solely to exploit structural properties of the original feasible schedules Sorted-F and Optimal (e.g., saturation/efficiency facts), rather than to claim feasibility of the intermediates.
We now begin the first step of the proof. Let $\{\mathcal{X}_{k}\}$ denote the sequence of batches iteratively generated in Phase 1 of Sorted-F. Let the requests in $\mathcal{X}_{k}$, ordered by increasing response length, be denoted as $r_{k,1},r_{k,2},\ldots,r_{k,n_{k}}$, where $n_{k}=|\mathcal{X}_{k}|$.
Definition 4.3 (Separation transformation $\textup{{Sorted-F}}\mapsto\textup{{Sorted-F}}_{\mathrm{separate}}$)
Fix the batch sequence $\{\mathcal{X}_{k}\}_{k\geq 1}$ produced by Phase 1 of Sorted-F. Let $p_{k,j}$ be the initiation time of request $r_{k,j}\in\mathcal{X}_{k}$ under Sorted-F. We construct $\textup{{Sorted-F}}_{\mathrm{separate}}$ by (i) enforcing no inter-batch overlap, (ii) enforcing simultaneous initiation within each batch, and (iii) making consecutive batches tight (i.e., no idle gap between them).
Define recursively:
$B_{1}:=\max_{1\leq j\leq n_{1}}p_{1,j},\qquad p^{\prime}_{1,j}:=B_{1},\ \ j=1,\ldots,n_{1},\qquad C^{\prime}_{1}:=\max_{1\leq j\leq n_{1}}\{p^{\prime}_{1,j}+o_{1,j}\}.$
For each $k\geq 2$, set
$B_{k}:=C^{\prime}_{k-1},\qquad p^{\prime}_{k,j}:=B_{k},\ \ j=1,\ldots,n_{k},\qquad C^{\prime}_{k}:=\max_{1\leq j\leq n_{k}}\{p^{\prime}_{k,j}+o_{k,j}\}.$
Thus all requests in $\mathcal{X}_{k}$ start simultaneously at time $B_{k}$, and batch $k$ starts exactly when batch k1k-1 finishes, i.e., $B_{k}=C^{\prime}_{k-1}$ for all $k\geq 2$.
The transformation in Definition 4.3 is schematically illustrated in Figure 2.
Figure 2: Schematic illustration of Definition 4.3 (from Sorted-F to $\textup{{Sorted-F}}_{\mathrm{separate}}$). The horizontal axis is time; each block represents a request whose width equals its response length.
@@
For any $\mathcal{I}$, the total end-to-end latency of Sorted-F is bounded above by that of $\textup{{Sorted-F}}_{\mathrm{separate}}$, i.e.,
@@
Proof 4.5
Let $p_{k,j}$ and $p^{\prime}_{k,j}$ denote the initiation times of request $r_{k,j}$ under Sorted-F and $\textup{{Sorted-F}}_{\mathrm{separate}}$, respectively. It suffices to show
$p_{k,j}\leq p^{\prime}_{k,j}\qquad\forall k\geq 1,\ \forall j\in\{1,\ldots,n_{k}\},$ (2)
since response lengths are unchanged.
We first show that Sorted-F has no idle gap between consecutive Phase 1 batches: for every $k\geq 2$,
$\max_{1\leq j\leq n_{k}}p_{k,j}\ \leq\ \max_{1\leq j\leq n_{k-1}}\bigl(p_{k-1,j}+o_{k-1,j}\bigr).$ (3)
Let $T_{k-1}:=\max_{1\leq j\leq n_{k-1}}(p_{k-1,j}+o_{k-1,j})$. If $\max_{j}p_{k,j}>T_{k-1}$, let $r_{k,\ell}$ be the first request in $\mathcal{X}_{k}$ (in the Phase 1 order) with $p_{k,\ell}>T_{k-1}$. Then $r_{k,\ell}$ is still pending at time $T_{k-1}$. However, at each timestep Phase 2 scans $\mathcal{I}^{\prime}$ in order and admits the longest feasible prefix; at time $T_{k-1}$ all requests in $\mathcal{X}_{k-1}$ have completed, which can only relax the feasibility checks. Since $\mathcal{X}_{k}$ is feasible in Phase 1 (Algorithm 1, lines 3–4) and feasibility is monotone under taking subsets, $r_{k,\ell}$ must be admissible no later than $T_{k-1}$, a contradiction. This proves (3).
We now verify (2). For $k=1$, Definition 4.3 sets $p^{\prime}_{1,j}=B_{1}=\max_{j^{\prime}}p_{1,j^{\prime}}\geq p_{1,j}$. For $k\geq 2$, Definition 4.3 sets $p^{\prime}_{k,j}=B_{k}=C^{\prime}_{k-1}$ for all $j$, where
$C^{\prime}_{k-1}=\max_{1\leq j\leq n_{k-1}}\{p^{\prime}_{k-1,j}+o_{k-1,j}\}.$
Since $p^{\prime}_{k-1,j}=B_{k-1}\geq p_{k-1,j}$, we have
$C^{\prime}_{k-1}\geq\max_{1\leq j\leq n_{k-1}}(p_{k-1,j}+o_{k-1,j})\geq\max_{1\leq j\leq n_{k}}p_{k,j},$
@@
$\sum_{r_{i}\in\mathcal{X}_{k}}(s_{i}+o_{i})\;>\;M-L(M),$
where $L(M):=\max_{i\in[n(M)]}\{s_{i}+o_{i}\}=\mathrm{o}(M)$ is the peak per-request memory defined in Section 2.
@@
Definition 4.8 (Grouping transformation $\textup{{Sorted-F}}_{\mathrm{separate}}\mapsto\textup{{Sorted-F}}_{\mathrm{group}}$)
Given an increasing index sequence $0=b_{0}<b_{1}<b_{2}<\cdots$, define grouped workloads
@@
where “++” means concurrent initiation (analytically) of all requests from the constituent batches. Formally, let $p^{\prime}_{k,j}$ be the initiation times in $\textup{{Sorted-F}}_{\mathrm{separate}}$. For each group $m$, set the group anchor time as $\tau_{m}:=\min_{1\leq j\leq n_{b_{m-1}+1}}p^{\prime}_{b_{m-1}+1,j}$, and for every batch $k\in\{b_{m-1}+1,\ldots,b_{m}\}$, shift all its requests by the same amount so that its earliest initiation aligns to $\tau_{m}$. The resulting initiation times define $\textup{{Sorted-F}}_{\mathrm{group}}$.
Figure 3 schematically illustrates Definition 4.8.
Figure 3: Schematic illustration of Definition 4.8 (from $\textup{{Sorted-F}}_{\mathrm{separate}}$ to $\textup{{Sorted-F}}_{\mathrm{group}}$).
Proof 4.9
Proof of Theorem 4.7: Each group $\mathcal{Y}_{m}$ is constructed such that $b_{m}$ is the smallest integer $k>b_{m-1}$ satisfying the inequality
@@
Next, we quantify how the processing times (i.e., the maximum response length within a batch) decay across batches inside each group. This is captured by the following lemma.
@@
The proof of Lemma 4.10 appears in Appendix 9. By Lemma 4.10, we obtain
$\sum_{i=1}^{a_{m}}o_{b_{m-1}+i,n_{b_{m-1}+i}}$ $\leq\sum_{i=1}^{a_{m}}\left(\frac{1}{2}\right)^{\lfloor\frac{a_{m}-i}{2}\rfloor}\cdot o_{b_{m},n_{b_{m}}}\leq 2\cdot\frac{1-\left(\frac{1}{2}\right)^{\lfloor\frac{a_{m}-1}{2}\rfloor}}{1-\frac{1}{2}}\cdot o_{b_{m},n_{b_{m}}}\leq 4\cdot o_{b_{m},n_{b_{m}}}.$ (4)
At this point, it remains to convert (4) into a TEL comparison. We rewrite $\mathrm{TEL}(\textup{{Sorted-F}}_{\mathrm{group}};\mathcal{I})$ and $\mathrm{TEL}(\textup{{Sorted-F}}_{\mathrm{separate}};\mathcal{I})$ in batch form, regroup by $\{\mathcal{Y}_{m}\}$, and apply (4) to bound each group’s contribution. The next lemma formalizes this bookkeeping step.
Lemma 4.11
For any $\mathcal{I}$,
$\mathrm{TEL}(\textup{{Sorted-F}}_{\mathrm{group}};\mathcal{I})>\frac{1}{4}\cdot\mathrm{TEL}(\textup{{Sorted-F}}_{\mathrm{separate}};\mathcal{I}).$
@@
Definition 4.12 (Alignment transformation $\textup{{Sorted-F}}_{\mathrm{group}}\mapsto\textup{{Sorted-F}}_{\mathrm{align}}$)
Consider the terminal batches $\{\mathcal{X}_{b_{m}}\}$ induced by the grouping in $\textup{{Sorted-F}}_{\mathrm{group}}$. For each terminal batch $\mathcal{X}_{b_{m}}$ with requests $\{r_{b_{m},1},\ldots,r_{b_{m},n_{b_{m}}}\}$, define its average response length
$\bar{o}_{b_{m}}=\frac{1}{n_{b_{m}}}\sum_{i=1}^{n_{b_{m}}}o_{b_{m},i}.$
The analytically aligned schedule $\textup{{Sorted-F}}_{\mathrm{align}}$ keeps the same initiation times as $\textup{{Sorted-F}}_{\mathrm{group}}$, but replaces every response length in each terminal batch $\mathcal{X}_{b_{m}}$ by $\bar{o}_{b_{m}}$ (all non-terminal batches remain unchanged).
Figure 4 schematically illustrates Definition 4.12.
Figure 4: Schematic illustration of Definition 4.12 (from $\textup{{Sorted-F}}_{\mathrm{group}}$ to $\textup{{Sorted-F}}_{\mathrm{align}}$).
@@
$\eta_{m}\;\geq\;\frac{\sum_{r_{i}\in\mathcal{Y}_{m}}(2s_{i}+o_{i})}{2M}\;>\;\frac{M-L(M)}{2M}\;\to\;\frac{1}{2}.$ (5)